40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3791     Entry time: Wed Oct 27 02:26:15 2010
Author: kiwamu 
Type: Update 
Category: IOO 
Subject: MC locked  

(Rana, Koji, Suresh, Yuta, Thanh, Kiwamu)

 MC was locked successfully !

 Instead of feeding back the signal to the MC length we just injected it to the NPRO pzt with a high voltage (HV) amplifier.

So now we can move on to an in-vac work which needs the main beam to align the stuff.

 


(mode matching to MC)

Suresh and Thanh (a visitor from ANU) improved the mode matching to the MC. 

As written in the entry #3779  the beam after the mode matching lenses were diverging.

It is supposed to converge from 1.7mm radius at the last lens to 1.6 mm radius at the middle point of MC1 and MC2.

They slided the last lens toward the MC to make it more collimated and roughly measured the beam size using a sensor card.

As a fine tuning, they looked at some higher order modes showing up in the MC2 camera, and tried reducing the higher order modes by slightly sliding the last lens.

(assuming the lens position doesn't so much change the alignment)

During the work we removed a steering mirror for green locking because it was on the way of the lens slider.

 

- - measured optics' distances - - 

25.5 cm  from 1st lens to the front surface of the EOM

5.5 cm  length of the EOM

24.5 cm from the front surface of the EOM to the 2nd lens (concave)

15.5 cm  from the 2nd lens to the 1st steering mirror in the zig-zag path

20.5 cm  from the steering mirror to the last lens

 

(preparation for locking) 

Rana, Yuta and Koji prepared an old instant amplifier which can produce +/-13V output instead of usual SR560s.

We added an offset (~5V) on the signal to make it within 0-10V which is the input range of the HV amplifier.

If we take SR560, it's probably not sufficiently wide range because they can handle handle only about +/-4V.

 

We strung a cable from Marconi via the RF stabilizer to the wideband EOM in order to drive the EOM at 24.5MHz.

SInce the EOM doesn't have 50 Ohm input impedance we had to put a 50 Ohm load just before the EOM in order to drive it efficiently.

From a medm screen we set the driving RF amplitude slider (C1:IOO_MCRF_AMPADJ) to 0.0, which provides the maximum RF power on it.

 

(locking mode cleaner)

At first we unlocked the PMC to see an offset in the error signal without any lights on the MC_REFL PD.

Then we adjusted the offset to zero on the MC servo screen.

At the beginning of the locking the PMC was not stable for some reasons during the MC was locked.

But after increasing the laser power to the MC twice bigger, it looks like the PMC and the MC are quite stable.

 

 

ELOG V3.1.3-